You are computing the future value of an ordinary annuity with annual payments of $650 each for four years. The annuity interest rate is 4 percent. The future value of each one of these four payments at the end of year 4 is:
Answer #1

[(1+r) -1 FV of Annuity = P P= Periodic Payment r=rate per period n = number of periods

100 1* 099 = AJ I - (100+1)

FV = 650 * 4.25

FV = $2,760.20